OpenStudy (anonymous):

A coin is tossed seven times. Find the probability of getting exactly four heads.

OpenStudy (raden):

exactly four heads means the events of 4H's (head) and 3T's (tail) the numbers events of 4H and 3T is = 7!/(4!3!) = .... and we knowed that the probability for 1H = 1/2 and also the probability for 1T = 1/2. assuming this event is independence for each tossed, so the total of probability of getting exactly four heads can be : P(4H, 3T) = 7!/(4!3!)*(1/2)*(1/2)*(1/2)*(1/2)*(1/2)*(1/2)*(1/2) = ... ?